OpenStudy (anonymous):

Input the equation of the given line in standard form. The line through (1, -4) and parallel to 2x + 3y = 4.

OpenStudy (anonymous):

2x + 3y = 4 3y = -2x + 4 y = -2/3x + 4/3 the slope of this line is -2/3. A parallel line will have the same slope. now we will use y = mx + b slope(m) = -2/3 (1,-4) x = 1 and y = -4 now we sub and solve for b -4 = -2/3(1) + b -4 = -2/3 + b -4 + 2/3 = b -12/3 + 2/3 = b -10/3 = b now we have : y = -2/3x - 10/3 -- add 2/3x to both sides 2/3x + y = -10/3 -- multiply by 3 2x + 3y = -10 ===> standard form